# Privacy as Social Norm: Systematically Reducing Dysfunctional Privacy Concerns on Social Media

### v2

@@ Abstract.
Privacy is essential to fully enjoying the benefits of social media. While fear around privacy risks can sometimes motivate privacy management, the negative impact of such fear, particularly when it is perceived as unaddressable (i.e., dysfunctional fear), can significantly harm teen well-being. In a co-design study with 136 participants aged 13-18, we explored how teens can protect their privacy without experiencing heightened fear. We identified seven different sources of dysfunctional fear, such as ‘fear of a hostile environment’ and ‘fear of overstepping privacy norms.’ We also evaluated ten designs, co-created with teen participants, that address these fears. Our findings suggest that social media platforms can mitigate dysfunctional fear without compromising privacy by creating a culture where privacy protection is the norm through default privacy-protective features. However, we also found that even the most effective privacy features are not likely to be adopted unless they balance the multifaceted and diverse needs of teens. Individual teens have different needs—for example, public and private account users have different needs—and teens often want to enjoy the benefits they get from slightly reducing privacy and widening their social reach. Given these considerations, augmenting default privacy features by allowing them to be toggled on and off will allow individual users to choose their own balance while still maintaining a privacy-focused norm.
social media, privacy, teens, adolescents
@@ 1. Introduction
In this paper, we contribute empirical evidence of dysfunctional fear in the context of teen privacy on social media. We also provide an overview of features and affordances that could effectively reduce both privacy concerns and risks. Furthermore, we identify moments where teens’ privacy needs collide with their peers’ privacy expectations and suggest methods for platforms to foster a culture where privacy is considered the norm. We hope this work will prompt designers and researchers to consider the adverse effects of excessive or unaddressed privacy concerns that teens may experience online and systematically empower adolescent users to protect themselves while fully reaping the benefits of social media.
@@ 2.3. Toward a Community-level Privacy Support for Teens on Social Media
As Emily Weinstein and Carrie James detail in their book (Weinstein and James, 2022), modern teenagers face unprecedented challenges in managing their online presence. The widespread practice of peers documenting and archiving digital interactions has created an environment of constant surveillance. Young people must navigate the threat of public shaming on social media, where past mistakes can be exposed and lead to widespread criticism or ostracism. The growing trend of saving digital evidence - whether through screenshots, recordings, or saved messages - has intensified these privacy concerns. This culture of perpetual documentation makes adolescents feel increasingly powerless over their personal information and digital reputation. The knowledge that any misstep could be preserved and shared creates significant anxiety among teens, who struggle to maintain autonomy over their online identities in this atmosphere of persistent observation and potential judgment. As such, the complexity and perceived lack of control often leads to “network defeatism” (De Wolf, 2020) among teens, a sense of fatalism regarding privacy decisions on social media (De Wolf, 2020; Hargittai and Marwick, 2016). This reduced sense of self-efficacy in privacy control hampers teens’ ability to address privacy issues and enhance their social media experiences (Lee et al., 2008; Jozani et al., 2020). Furthermore, teens are more likely to overlook their vulnerability when they feel less efficacious, leading to disengagement from privacy protection (Chou and Chou, 2023). Adults’ well-meaning warnings about the permanence of online actions often add to this disempowerment, instilling a sense of fear and cynicism (Weinstein and James, 2022).
@@ 3.1. Co-design Interviews
During exit interviews, we asked participants about the barriers (if any) that prevent them from comfortably sharing on social media. To do so, we used the Miro platform (mir, [n. d.]) to create a virtual whiteboard with nine sticky notes, each representing the improvements they seek on social media (and its related subcategories of concerns) identified in our survey. Examples of the larger themes on changes sought were “prevent posts from being shared out of context,” while sub-themes of privacy concerns included “Content being shared with strangers” and “Screenshots taken / content saved.” (The full list of the sticky notes is available in the supplementary materials.) We used these to prompt a conversation where participants highlighted the barriers most relevant to them.
@@
In the final phase, the primary author coded all transcripts to refine the codebook. The two co-authors then coded 20 of the 38 transcripts, validating the clarity and validity of the codebook and the assigned codes. Throughout each phase, the team convened to discuss higher-level themes, disagreements, or ambiguities related to the codebook. Higher-level themes included “Fear that teens experience over privacy,” “Existing strategies for mitigating fear,” “Designs that affect teens’ fear,” and “Implications of the designs.” Sub-codes included “Ripple audience,” “Perspective taking,” “Prompts for self-regulation,” and “Privacy-utility trade-off.” Each iteration of the codebook, complete with definitions and example quotes for each code, is available in the supplementary materials.
@@ 3.2. Design Evaluation Survey
Materials and Procedures. For each of the ten design idea themes developed during the inductive coding process of the co-design interviews, the first and fourth authors—both advanced-degree students in technology design programs—created high-fidelity prototypes of ten sample features (See Figure 1 for sample prototypes and the supplementary materials for the entire set of prototypes) that were provided to the survey respondents. These authors selected specific implementations for each design based on their potential to probe the defining characteristics of each design idea.
@@
We hypothesized that the effectiveness of, and thus the reactions to, the features would significantly differ between public and private account owners. We also anticipated they would have different types and extents of privacy concerns. Consequently, we created two versions of the same survey, one for public and another for private account owners. Each version asked participants to consider the features and privacy concerns in the context of the public or private account they most frequently use. We permitted participants to respond to both surveys if they owned both types of accounts. The complete wording of the two versions of surveys can be found in the supplementary materials.
@@
Each response was reviewed, and we excluded those that appeared to be 1) submitted multiple times by an individual, 2) generated by generative AI, or 3) not genuine, such as responses not addressing the question or only offering vague positive comments such as “I would love this” without any evidence of having reviewed the prototypes throughout the entire survey. The detailed process and criteria we employed for this process are available in the supplementary materials. After this filtering process, data from 72 private account owners and 64 public account owners remained. Among these, 18 participants responded to both surveys. The demographics of the 118 participants (who provided 136 responses) are detailed in Appendix A.
@@ 5.2. A Design Agenda for Privacy Assurance
Cultivate a Culture of Privacy. Teens often experience dysfunctional fear of privacy issues on social media, leading to a perceived lack of self-efficacy. Paradoxically, this can result in them surrendering their privacy, believing issues to be inevitable. This suggests that the current model of individual user control is insufficient, highlighting the need for platforms to provide both protection and reassurance to effectively address teens’ privacy concerns.
@@
Recognize Multifaceted Needs. Based on survey respondents’ concerns, we identified four main areas of potential pitfalls of the privacy features: negative impact on other users, restriction of social media utility, user burden (e.g., features triggered without active engagement), and community-dependency (where individual engagement alone is insufficient to reduce privacy risks). We analyzed each prototype against these criteria, summarizing the results in Table 2. As anticipated, features without these drawbacks—Personal Information Alert and Red Flag & Follow-up—were most positively received. Conversely, Auto-delete (with save), which exhibited all four issues, was the least favored.
@@
Support Diverse Needs. Our study revealed that teens have varied privacy requirements, particularly evident in the contrast between private and public account users. While private account holders prioritize strict privacy controls, public account users, such as aspiring artists, seek a balance between visibility and protection. To address this spectrum of needs, participants proposed design improvements that offer nuanced privacy options. For public accounts, suggestions included disconnecting personal information from public visibility, systematically filtering out users who fail to meet community standards, and enhancing interaction transparency through better-organized viewer lists. These features aim to enable public account holders to engage with a larger audience without fully compromising their privacy, demonstrating the need for flexible privacy solutions that cater to diverse user goals.
@@ 6. Conclusion
In this paper, we explored the privacy-related fears experienced by teens on social media and investigated how design can mitigate these fears without compromising privacy protection. Our study involved co-design interviews with 19 teens and a design evaluation survey with 136 U.S. teen participants. Our findings indicate that teens frequently experience vague and persistent fears about privacy that they feel unable to manage effectively, leading to dysfunctional fear that inhibits their use of social media and negatively impacts their well-being. These fears fall into three primary categories: fear of uncontrolled audience reach, fear of online hostility, and fear of personal privacy missteps. Teens suggested various design improvements to address these fears, such as prompts for self-regulation and clearer communication of privacy norms. We developed ten concrete prototypes for these design ideas, some of them based on existing features, with an emphasis on increasing user agency over the status quo. Survey participants found these design prototypes helpful in reducing their privacy concerns and enhancing their sense of control on social media. Our study provides empirical evidence of dysfunctional fear in teens’ privacy management on social media and offers guidelines to support user empowerment. By incorporating system-wide guidance, enforcement, and options to make privacy both the norm and a priority, social media platforms can help teens navigate complex social interactions while maintaining privacy.
@@ Appendix A Demographics Data
Table 2. provides demographic information about the participants involved in the co-design study. It includes gender distribution, average age, and racial background of the participants. It describes the platforms that participants share content on frequently and comfortably. There are 2 columns with the categories in the first column on a grey background with black text, and the data in the second column on a white background with black text. Gender identity Girls (63%), Boys (26%), Non-binary or third gender (5%), Boy and non-binary or third gender (5%) Age red13 (10.5%), 14 (10.5%), 15 (15.8%), 16 (21.1%), 17 (21.1%), 18 (21.1%) Race White (53%), Asian or Asian-American (32%), Black or African American (11%), White and Asian or Asian-American (5%) Hispanic or Latin-American Origin No (84%), Yes (16%) Social Media Usage Instagram (100%), BeReal (84%), Snapchat (74%), Twitter (68%), TikTok (63%), Discord (11%), Reddit (11%), Tumblr (5%) Platforms Where Participants Reported Sharing Most Frequently Instagram (37%), BeReal (32%), Snapchat (11%)), Twitter (5%), YouTube (5%), Discord (5%), Rinsta (5%) Platforms Where Participants Reported Sharing Most Comfortably Instagram (26%), BeReal (26%), Snapchat (16%), Discord (16%), Rinsta (5%), Pinterest (5%), Finsta (5%)

### v3

@@ Abstract.
Through co-design interviews ($N=19$) and a design evaluation survey ($N=136$) with U.S. teens ages 13-18, we investigated teens’ privacy management on social media. Our study revealed that 28% of teens with public accounts and 15% with private accounts experience dysfunctional fear, that is, fear that diminishes their quality of life or paralyzes them from taking necessary precautions. These fears fall into three categories: fear of uncontrolled audience reach, fear of online hostility, and fear of personal privacy missteps. While current approaches often emphasize individual vigilance and restrictive measures, our findings show this can paradoxically lead teens to either withdraw from beneficial social interactions or resign themselves to accept privacy violations, viewing them as inevitable. Drawing on teen input, we developed and evaluated ten design prototypes that emphasize empowerment over fear, system-wide explicit emphasis on privacy, clear privacy norms, and flexible controls. Survey results indicate teens perceive these approaches as effectively reducing privacy concerns while preserving social benefits. Our findings suggest that platforms will be more likely to protect teens’ privacy and less likely to manufacture unnecessary fear if they include designs that minimize the impact on other users, have low trade-offs with existing features, require minimal user effort, and function independently of community behavior. Such designs include: 1) alerting users about potentially unintentional personal information disclosure and 2) following up on user reports.
social media, privacy, fear, teens, adolescents, social norm
@@ 1. Introduction
In this paper, we contribute empirical evidence of dysfunctional fear in the context of teen privacy on social media. We also provide an overview of features and affordances that could effectively reduce both privacy concerns and risks. Furthermore, we identify moments where teens’ privacy needs collide with their peers’ privacy expectations and suggest methods for platforms to foster a culture where privacy is considered the norm. We hope this work will prompt designers and researchers to consider the adverse effects of alarmism or prevention-focused strategies and, instead, systematically empower adolescent users to protect themselves while fully reaping the benefits of social media.
@@ 2.3. Toward a Community-level Privacy Support for Teens on Social Media
As Emily Weinstein and Carrie James detail in their book (Weinstein and James, 2022), modern teenagers face unprecedented challenges in managing their online presence. The widespread practice of peers documenting and archiving digital interactions has created an environment of constant surveillance. Young people must navigate the threat of public shaming on social media, where past mistakes can be exposed and lead to widespread criticism or ostracism. The growing trend of saving digital evidence—whether through screenshots, recordings, or saved messages—has intensified these privacy concerns. This culture of perpetual documentation makes adolescents feel increasingly powerless over their personal information and digital reputation. The knowledge that any misstep could be preserved and shared creates significant anxiety among teens, who struggle to maintain autonomy over their online identities in this atmosphere of persistent observation and potential judgment. As such, the complexity and perceived lack of control often leads to “network defeatism” (De Wolf, 2020) among teens, a sense of fatalism regarding privacy decisions on social media (De Wolf, 2020; Hargittai and Marwick, 2016). This reduced sense of self-efficacy in privacy control hampers teens’ ability to address privacy issues and enhance their social media experiences (Lee et al., 2008; Jozani et al., 2020). Furthermore, teens are more likely to overlook their vulnerability when they feel less efficacious, leading to disengagement from privacy protection (Chou and Chou, 2023). Adults’ well-meaning warnings about the permanence of online actions often add to this disempowerment, instilling a sense of fear and cynicism (Weinstein and James, 2022).
@@ 3.1. Co-design Interviews
During exit interviews, we asked participants about the barriers (if any) that prevent them from comfortably sharing on social media. To do so, we used the Miro platform (mir, [n. d.]) to create a virtual whiteboard with nine sticky notes, each representing the improvements they seek on social media (and its related subcategories of concerns) identified in our survey. Examples of the larger themes on changes sought were “prevent posts from being shared out of context,” while sub-themes of privacy concerns included “Content being shared with strangers” and “Screenshots taken / content saved.” (The full list of the sticky notes is available in the exit interview template in https://osf.io/8t7cy/) We used these to prompt a conversation where participants highlighted the barriers most relevant to them.
@@
In the final phase, the primary author coded all transcripts to refine the codebook. The two co-authors then coded 20 of the 38 transcripts, validating the clarity and validity of the codebook and the assigned codes. Throughout each phase, the team convened to discuss higher-level themes, disagreements, or ambiguities related to the codebook. Higher-level themes included “Fear that teens experience over privacy,” “Existing strategies for mitigating fear,” “Designs that affect teens’ fear,” and “Implications of the designs.” Sub-codes included “Ripple audience,” “Perspective taking,” “Prompts for self-regulation,” and “Privacy-utility trade-off.” Each iteration of the codebook, complete with definitions and example quotes for each code, is available in https://osf.io/8t7cy/).
@@ 3.2. Design Evaluation Survey
Materials and Procedures. For each of the ten design idea themes developed during the inductive coding process of the co-design interviews, the first and fourth authors—both advanced-degree students in technology design programs—created high-fidelity prototypes of ten sample features (See Figure 1 for sample prototypes and https://osf.io/8t7cy/) for the entire set of prototypes) that were provided to the survey respondents. These authors selected specific implementations for each design based on their potential to probe the defining characteristics of each design idea.
@@
We hypothesized that the effectiveness of, and thus the reactions to, the features would significantly differ between public and private account owners. We also anticipated they would have different types and extents of privacy concerns. Consequently, we created two versions of the same survey, one for public and another for private account owners. Each version asked participants to consider the features and privacy concerns in the context of the public or private account they most frequently use. We permitted participants to respond to both surveys if they owned both types of accounts. The complete wording of the two versions of surveys can be found in https://osf.io/8t7cy/).
@@
Each response was reviewed, and we excluded those that appeared to be 1) submitted multiple times by an individual, 2) generated by generative AI, or 3) not genuine, such as responses not addressing the question or only offering vague positive comments such as “I would love this” without any evidence of having reviewed the prototypes throughout the entire survey. The detailed process and criteria we employed for this process are available in https://osf.io/8t7cy/). After this filtering process, data from 72 private account owners and 64 public account owners remained. Among these, 18 participants responded to both surveys. The demographics of the 118 participants (who provided 136 responses) are detailed in Appendix A.
@@ 5.2. A Design Agenda for Privacy Assurance
From Privacy Paralysis to Platform Responsibility. Teens often experience dysfunctional fear of privacy issues on social media, leading to a perceived lack of self-efficacy. Paradoxically, this can result in them surrendering their privacy, believing issues to be inevitable. This suggests that the current model of individual user control is insufficient, highlighting the need for platforms to provide both protection and reassurance to effectively address teens’ privacy concerns.
@@
Mapping Privacy-Utility Tensions: A Framework for Feature Evaluation. Based on survey respondents’ concerns, we identified four main areas of potential pitfalls of the privacy features: negative impact on other users, restriction of social media utility, user burden (e.g., features triggered without active engagement), and community-dependency (where individual engagement alone is insufficient to reduce privacy risks). We analyzed each prototype against these criteria, summarizing the results in Table 2. As anticipated, features without these drawbacks—Personal Information Alert and Red Flag & Follow-up—were most positively received. Conversely, Auto-delete (with save), which exhibited all four issues, was the least favored.
@@
Beyond Binary Privacy: Public Accounts’ Privacy Protection Needs. Our study revealed that teens have varied privacy requirements, particularly evident in the contrast between private and public account users. While private account holders prioritize strict privacy controls, public account users, such as aspiring artists, seek a balance between visibility and protection. However, public account owners also reported twice as much experience with a dysfunctional fear of privacy risks than private account owners.
To address this spectrum of needs, participants proposed design improvements that offer nuanced privacy options. For public accounts, suggestions included disconnecting personal information from public visibility, systematically filtering out users who fail to meet community standards, and enhancing interaction transparency through better-organized viewer lists. These features aim to enable public account holders to engage with a larger audience without fully compromising their privacy, demonstrating the need for flexible privacy solutions that cater to diverse user goals.
@@ 6. Conclusion
In this paper, we explored the privacy-related fears experienced by teens on social media and investigated how design can mitigate these fears without compromising privacy protection. Our study involved co-design interviews with 19 teens and a design evaluation survey with 136 U.S. teen participants. Our findings indicate that teens frequently experience vague and persistent fears about privacy that they feel unable to manage effectively, leading to dysfunctional fear that inhibits their use of social media and negatively impacts their well-being. These fears fall into three categories: fear of uncontrolled audience reach, fear of online hostility, and fear of personal privacy missteps.
By incorporating system-wide guidance, enforcement, and options to make privacy both the norm and a priority, social media platforms can help teens navigate complex social interactions while maintaining privacy. Our study provides empirical evidence of dysfunctional fear in teens’ privacy management on social media and offers guidelines to support user empowerment. This work highlights the critical need to move beyond simply warning teens about privacy risks—which can lead to counterproductive anxiety and withdrawal—toward creating environments where privacy-conscious behavior is normalized and supported through thoughtful design.
Our findings emphasize that effective privacy protection for teens requires balancing safety with social connection, implementing proactive rather than fear-based approaches, and fostering community norms that support privacy-conscious behavior. Future research should continue to explore how platform design can help transform vague privacy anxieties into constructive privacy practices, enabling teens to fully participate in online social spaces while maintaining appropriate boundaries.
Acknowledgements.
The authors would like to acknowledge the CERES network, which provided support for this work. We additionally thank the anonymous reviewers for their detailed feedback and the participants for sharing their thoughts. We truly appreciate all the help. Alexis Hiniker is a special government employee for the Federal Trade Commission. The content expressed in this manuscript does not reflect the views of the Commission or any of the Commissioners.
@@ Appendix A Demographics Data
Table 2. provides demographic information about the participants involved in the co-design study. It includes gender distribution, average age, and racial background of the participants. It describes the platforms that participants share content on frequently and comfortably. There are 2 columns with the categories in the first column on a grey background with black text, and the data in the second column on a white background with black text. Gender identity Girls (63%), Boys (26%), Non-binary or third gender (5%), Boy and non-binary or third gender (5%) Age 13 (10.5%), 14 (10.5%), 15 (15.8%), 16 (21.1%), 17 (21.1%), 18 (21.1%) Race White (53%), Asian or Asian-American (32%), Black or African American (11%), White and Asian or Asian-American (5%) Hispanic or Latin-American Origin No (84%), Yes (16%) Social Media Usage Instagram (100%), BeReal (84%), Snapchat (74%), Twitter (68%), TikTok (63%), Discord (11%), Reddit (11%), Tumblr (5%) Platforms Where Participants Reported Sharing Most Frequently Instagram (37%), BeReal (32%), Snapchat (11%)), Twitter (5%), YouTube (5%), Discord (5%), Rinsta (5%) Platforms Where Participants Reported Sharing Most Comfortably Instagram (26%), BeReal (26%), Snapchat (16%), Discord (16%), Rinsta (5%), Pinterest (5%), Finsta (5%)